OpenStudy (anonymous):

What is the distance between the points (-1, 5) and (3, 7)? Round your answer to the nearest hundredth.

OpenStudy (anonymous):

\(Distance = \sqrt{\Delta x^2 + \Delta y^2}\) \(\Delta x=x_2-x_1\) \(\Delta y=y_2-y_1\) Go.

OpenStudy (anonymous):

\[d = \sqrt{(x _{2}-x _{1})^{2}+ (y _{2}-y _{1})^{2}} = \sqrt{(3+1)^{2}+(7-5)^{2}} = \sqrt{20} = 2\sqrt{5}\]